OpenStudy (anonymous):

what is the gravitational force between two masses if one is 75kg and the other 54kg and their centers are seperated by a distance of .45 meters

OpenStudy (anonymous):

do you want the formula?

OpenStudy (anonymous):

newton s law of gravitation states that F = GMm/R^2 g is gravitational constant =6.67*10^(-11) Nm^2Kg^(-2) R Is the distance between them. M & m are the masses of two bodies

OpenStudy (anonymous):

i hope it will help you.

OpenStudy (anonymous):

thanks it will I appreciate it hopefully I can calculate it right

OpenStudy (anonymous):

welcome.